# Information-Driven Design of Imaging Systems

### v4

@@ Acknowledgements
What is the volume of the space of possible signals? Given that all signals have energy $\leq 1$, the vector that defines their representation in energy coordinates must have $L_{1}$ norm also $\leq 1$. Thus, all signals must correspond to points that lie inside the positive orthant of the $L_{1}$ unit ball. However, not every point in this space will correspond to a valid signal: for example, a vector with a single $1$ and the rest 00s will not be possible, because this would entail concentrating all of the signal’s energy within a single pixel.
It is unclear to us if there is an analytical expression that defines the set of possible signals, but we can investigate the size of this set empirically. To do so, we set up an optimization problem in which we pick a fixed object and a target energy coordinate representation of a signal (e.g. a vector with a single $1$ and the rest 00s). We then find the optimal point spread function that brings the object closest to the target signal by optimizing an encoder to minimize this distance using gradient descent. Repeating this experiment over a grid of target signals shows which signals can be reached, and which cannot, thereby revealing the limits of the space of possible signals.
Repeating this experiment with different fixed objects illustrates an important insight about encoders: Their range is object dependent (Fig. S5). This is a direct result of their physical constraints. The 1D encoder in this simulation is representative of imaging systems governed by intensity point spread functions [93]. Such encoders have at least two important physical constraints: 1) they can only reduce energy (if they attenuate light) or preserve it. They cannot, for example, encode a dim object to a signal with greater energy. 2) They can only disperse, and not re-concentrate, energy. Every point spread function (under the constraints of non-negativity and linear shift invariance) can only map objects to signals that are blurrier versions of themselves.
@@
To find the best achievable distribution, we first use Information-Driven Encoder Analysis Learning (IDEAL) (See Information-Driven Encoder Analysis Learning (IDEAL)) to learn the optimal encoder. We then generate random objects and encode them with this optimal encoder to sample from the distribution of encoded signals, add noise, and estimate the mutual information. We use the PixelCNN-based estimator [^2] We used PixelCNN for convenience, though estimators designed for 1D signals [94, 95] might provide better accuracy., treating the signals as 2D images to match the estimator’s design.
@@
Next, we tested the effect of signal bandwidth on information capacity. Optical imaging systems are often characterized in terms of their space-bandwidth product [96], and the space-bandwidth product is often used synonymously with the word “information.” A more accurate term for the space-bandwidth product is “degrees of freedom,” since it quantifies the potential complexity of an electromagnetic field wave propagating through the system [90]. Information (in Shannon’s entropy/mutual information sense) also depends on the signal-to-noise ratio and the object-dependent ability of encoders to map to distinct signals. In our numerical simulation, the spatial extent of signals is fixed, so increasing the bandwidth of the signal increases the space-bandwidth product. We found that captured information increases linearly with the space-bandwidth product, with rate of increase depending on the object distribution (Fig. S7b, center).
@@
Finally, we examined the effect of sampling density on the amount of information that can be encoded. For a fixed bandwidth, the sampling density determines the number of pixels in the signal. The Nyquist sampling theorem [97, 89] states that a bandlimited signal can be perfectly reconstructed from its samples if the sampling density is at least twice the bandwidth. However, in the presence of noise, even when sampling at the Nyquist rate, there remains residual uncertainty about the signal, and additional samples can reduce this uncertainty. Experimentally, we found that, as expected, increasing the sampling density increases the amount of information, even beyond the Nyquist rate (Fig. S7b, right). However, the additional increases in information were progressively smaller. These results were consistent across both different object distributions and different average signal-to-noise ratios (for a delta function object distribution).
@@
$\delta(\mathbf{o}-\mathbf{o}_{1})$ is $1$ when the object is $\mathbf{o}_{1}$ and 00 otherwise, and similarly for $\delta(\mathbf{o}-\mathbf{o}_{2})$.
@@
Since the goal of the imaging system in this simple example is to classify the object as being a single point source or two point sources [91], the decoder is a binary classifier that takes in the noisy measurement and outputs a decision as to whether the object was a single point or two points.
@@
The minimal model from the previous sections makes simplifying assumptions and abstracts away details that are relevant to real imaging systems. In the section we generalize to a more complex model, and explain the deeper insights and capabilities such a model can provide.
@@
S1.5.1 Encoder uncertainty and calibration
In the minimal model in the previous sections, the encoder function is fixed, so all variations in the noiseless images come from variations in the objects. This enables the simplifying assumption that the mutual information between the noisy measurements and noiseless images equals the mutual information between the noisy measurements and the objects: $I(\mathbf{Y};\mathbf{X})=I(\mathbf{Y};\mathbf{O})$. As a result, estimating the former also reveals how much information about the objects is contained in the noisy measurements.
However, in practice the encoder may vary for different objects. For example, dust or imperfections in optical components, fluctuations in illumination brightness, and non-uniform resolution across the field of view can all change from measurement to measurement. Imaging systems are generally designed to minimize such variations - lenses provide uniform resolution across the field of view, light sources are engineered for stability, and care is taken to avoid dust in the optical path. But some degree of variation is inevitable.
As a result, observing noisy measurements provides information both about the object and about the measurement system itself. Not accounting for encoder variations will lead to overestimates of the object information contained in the measurements.
The randomness of encoders is qualitatively different than the randomness caused by detection noise. Detection noise arises from inherently unpredictable quantum phenomena like photon arrival times. In contrast, the randomness of the encoder models a fixed but unknown distortion applied to the signal that, if learned, can be accounted for and will no longer be a source of randomness.
Mathematically, the encoder sits outside the Markov chain from object to image. Thus the data processing inequality does not imply permanent information loss as a result of its randomness. Through calibration, the nature of the distortion can be learned. Once characterized, the apparent randomness in future measurements is reduced, enabling recovery of information about the object. There are many examples of this concept in practice - deconvolution algorithms can incorporate spatially varying point spread functions, shading corrections can be applied to account for varied illumination of the field of view [98, 99], etc.
A similar calibration approach can be taken with information estimators to avoid over-estimating the object information in the measurement. We describe a theoretical framework for implementing this in Section S2.7, and demonstrate its practical importance through experiments with LED array microscopy in Section S4.1. These experiments show how failing to account for encoder uncertainty can significantly impact information estimates, particularly in sensitive imaging systems.
S1.5.2 Task-specific information
Another useful generalization is to explicitly model the relationship between the decoding task and the object being imaged [23, 68]. For example, a perfect decoder for a specific task may require only a subset of the information needed to fully describe the object’s physical properties. Alternatively, even complete knowledge of the object’s physical properties may be insufficient to perfectly predict the task outcome - no encoder could encode enough task-relevant information to enable a flawless decoder.
To account for this, the model includes an additional random variable:
$\boldsymbol{\mathrm{T^{*}}}$ the true value of the decoding task, which sits upstream of the object $\mathbf{O}$ in the Markov chain. This expresses that the object may not contain all task-relevant information, and also may contain additional irrelevant details.
The full Markov chain is now $\boldsymbol{\mathrm{T^{*}}}\rightarrow\mathbf{O}\rightarrow\mathbf{X}\rightarrow\mathbf{Y}$. The key quantity of interest is $I(\boldsymbol{\mathrm{T^{*}}};\mathbf{Y})$, the mutual information between the true task value and the noisy measurement. This quantifies how much task-relevant information makes it through the entire imaging system. It can be substantially less than $I(\mathbf{O};\mathbf{Y})$, as it excludes irrelevant object details.
The greater the disparity between $I(\boldsymbol{\mathrm{T^{*}}};\mathbf{Y})$ and $I(\mathbf{O};\mathbf{Y})$, the less predictive estimated information may be of decoder performance.
This seems especially likely when the task is very simple, capturing just a small fraction of the information in the measurements about the full object state. As an example, in Section S4.2 we observe a scenario where a simple classification task requires only a small fraction of the information present in the full measurements.
@@
As described in Formulation of information estimator, our probabilistic model represents noiseless images as a random 2D array $\mathbf{X}$ with distribution $p(\mathbf{x})$, where each element quantifies the energy at a single pixel. The corresponding noisy measurements form array $\mathbf{Y}$ with distribution $p(\mathbf{y})$. These distributions determine how much object-relevant information survives the measurement process.
@@
While imaging system performance typically varies across the field of view, we can quantify overall information content using a single value: the mutual information rate (information per pixel). This parallels how optical resolution is often characterized by a single number despite spatial variations [93].
@@
Mutual information estimation is a well-studied problem in many fields, including machine learning, neuroscience, and computational biology, and a number of different approaches have been proposed, some of which attempt to estimate mutual information directly [33, 34, 35], and others which try to infer its value by computing upper or lower bounds [43, 36]. It is a challenging statistical problem to solve in general, and many estimators and bounds are known to suffer from high bias and/or variance, particularly on high-dimensional problems [43, 36].
@@
Like mutual information estimation, there is a large body of literature on entropy estimation and many different approaches have been proposed [100], including those that form an estimate of the probability density function $\hat{p}(\mathbf{y})$ and plug it into the definition of entropy and those that estimate entropy based on the similarity of samples from $p(\mathbf{y})$ [101, 102, 103]. Both approaches, however, face difficulties in high dimensions.
@@
(A technical note: in optical systems, the outcomes of the random variables $\mathbf{X}$ and $\mathbf{Y}$ in our model are discrete, because they are 2D arrays of pixels, where each pixel takes on an intensity value that is an integer number of photons. However, for computational simplicity, we will approximate them as continuous random variables and use differential entropy instead of discrete entropy [72, 69]. These approximations break down at very low photon counts, less than $\sim 20$ photons, so in this paper, we use only data with photon counts greater than this. A possible direction of future work is to extend our approach to work with discrete random variables, which would allow it to be applied to data with lower photon counts.)
@@
$p(\mathbf{y}\mid\mathbf{x})$ embodies the various sources of noise in the detection process, including photon shot noise, detector read noise, etc. Here, we utilize established analytic models of detection noise in optical imaging [32]. Empirical results suggest that the true noise in systems deviates from these models in low-light conditions [104], but since our experiments are conducted in the high-light regime, we will assume that these models are accurate for the purposes of this paper. A possible direction for future work is to learn more accurate noise models from the data, as was done in [105].
@@
The entropy of a (scalar) Gaussian distribution $\mathcal{N}(\mu,\sigma)$ is [72]:
@@
Plugging this result into Equations 15 and 11 yields:
@@
Images with high photon counts are fundamentally limited by shot noise - randomness in photon arrival times due to the quantum nature of light. This shot noise follows a Poisson distribution with rate parameter equal to the expected number of photons at each pixel. When shot noise is the dominant source of noise, it can be accurately approximated by a Gaussian distribution with equal mean and variance [32].
@@
We estimate this entropy by computing an upper bound, an approach that has proven more accurate than alternative bounds in high-dimensional settings [43]. Our method fits a parametric model $p_{\mathbf{\theta}}(\mathbf{y})$ to the empirical distribution of noisy images using maximum likelihood estimation. The optimal parameters $\hat{\theta}_{\mathrm{MLE}}$ are found by minimizing the negative log likelihood of the observed data:
@@
This loss function, $-\mathbb{E}\left[\log p_{\theta}(\mathbf{Y})\right]$, is also known as the cross-entropy between the model distribution $p_{\theta}(\mathbf{y})$ and the empirical distribution $p(\mathbf{y})$.
@@
Other models could be used within this framework, provided they allow direct evaluation of likelihood functions. Recent neural network architectures such as transformers [76], normalizing flows [106], and diffusion models [107, 108] offer promising directions for future work.
While generative models are often evaluated by the visual quality of their samples, this metric does not necessarily correlate with cross-entropy performance [109] - though some state-of-the-art models can achieve both [110]. We selected PixelCNN for its demonstrated effectiveness in minimizing test set cross-entropy [47], combined with its relative simplicity and computational efficiency. However, the rapid advancement of generative models suggests that more efficient architectures providing tighter entropy bounds for a given computational efficiency are likely to emerge.
@@
To reduce model complexity and quantify performance over the full field of view with a single scalar quantity, we can make the simplifying assumption that images are stationary stochastic processes. This is the probabilistic analog of assuming constant optical resolution across the field of view [93].
@@
Stationary processes have another useful property: a constant entropy per pixel, called the entropy rate [72] (explained in Figure 8 of [87]). This rate can be defined in two equivalent ways:
@@
Our most flexible model is the PixelCNN [45, 46, 47], which uses neural networks to capture complex pixel dependencies through an autoregressive approach.
@@
This factorization does not require specific assumptions about the joint distribution, so in theory it can be applied to any distribution. Creating the full model requires $D$ conditional distribution models, each taking between 1 and $D-1$ previous pixel values as input and outputting a 1D probability distribution for the next pixel. While this could be implemented with $D$ separate models, PixelCNN uses masked convolutions in a single neural network to model all conditionals simultaneously, dramatically reducing computational complexity.
Following previous work [45, 46], our architecture uses a series of masked convolutions to maintain the autoregressive ordering, producing a conditional probability distribution at each pixel.
However, we make one crucial modification: instead of the standard softmax output layer, we use a mixture of Gaussians parameterized by the network (a mixture density network [111]). This strategy is capable in theory of approximating any conditional probability distribution, in the same way that a neural network is capable of approximating any function [112]. This modification is essential for estimating entropy in our framework, because it means the output distributions will be continuous probability densities, instead of discrete probability mass functions. Since our noise models $p(\mathbf{y}\mid\mathbf{x})$ are continuous (Gaussian), our entropy estimates for $H(\mathbf{Y})$ must also use continuous distributions to allow proper subtraction of the conditional entropy $H(\mathbf{Y}\mid\mathbf{X})$.
@@
While our models’ primary purpose is likelihood evaluation for entropy estimation (Section S2.4), generating samples provides valuable insight into model behavior. For simple models like the stationary Gaussian process, sample quality correlates well with likelihood accuracy, making visual inspection a useful diagnostic tool. However, this correlation breaks down for more complex models like PixelCNN [109], though samples remain helpful for understanding model behavior.
@@
Computing the likelihood over an image larger than the length $D$ model the pixel was trained on can be accomplished by adding additional terms to this sum. For example, computing the log likelihood of an $D+1$ length image would require adding an additional term to the sum of the log likelihood of the the $D+1$ pixel conditioned on the previous $D-1$ pixels (the maximum extent of the model) that preceded it:
@@
Details of the likelihood computations for PixelCNNs are described in [45]. For stationary Gaussian processes, there is a closed form solution for finding the mean $\mu_{D}$ and variance $\sigma^{2}_{D}$ of a 1-dimensional Gaussian distribution conditioned on $D-1$ previous values. This involves decomposing the covariance matrix into a top left $(D-1)\times(D-1)$ block $\mathbf{\Sigma}_{1,1}$, a top right $(D-1)\times 1$ column vector $\mathbf{\Sigma}_{1,2}$, and a bottom left $1\times(D-1)$ row vector $\mathbf{\Sigma}_{2,1}$:
[…]
Given $D-1$ previous values $\mathbf{y}_{D-1}$, the mean and variance of the $D$th value can be computed as:
@@
Similarly, sampling images larger than the patch size on which the models were trained can be accomplished by iteratively sampling each pixel conditioned on the previous $D-1$ pixels.
S2.7 Accounting for encoder uncertainty
Calibration presents a fundamental challenge in applying information estimation to experimental imaging systems (Section S1.5.1). While the basic approach treats object properties and detection noise as the primary sources of measurement randomness, real systems introduce additional variations through optical imperfections, mechanical instabilities, and illumination fluctuations. These variations can inflate information estimates by misattributing system randomness to object information.

### v5

@@ Acknowledgements
What is the volume of the space of possible signals? Given that all signals have energy $\leq 1$, the vector that defines their representation in energy coordinates must have $L_{1}$ norm also $\leq 1$. Thus, all signals must correspond to points that lie inside the positive orthant of the $L_{1}$ unit ball. However, not every point in this space will correspond to a valid signal: for example, a vector with a single $1$ and the rest 0s will not be possible, because this would entail concentrating all of the signal’s energy within a single pixel.
It is unclear to us if there is an analytical expression that defines the set of possible signals, but we can investigate the size of this set empirically. To do so, we set up an optimization problem in which we pick a fixed object and a target energy coordinate representation of a signal (e.g. a vector with a single $1$ and the rest 0s). We then find the optimal point spread function that brings the object closest to the target signal by optimizing an encoder to minimize this distance using gradient descent. Repeating this experiment over a grid of target signals shows which signals can be reached, and which cannot, thereby revealing the limits of the space of possible signals.
Repeating this experiment with different fixed objects illustrates an important insight about encoders: Their range is object dependent (Fig. S5). This is a direct result of their physical constraints. The 1D encoder in this simulation is representative of imaging systems governed by intensity point spread functions [goodmanIntroductionFourierOptics2005]. Such encoders have at least two important physical constraints: 1) they can only reduce energy (if they attenuate light) or preserve it. They cannot, for example, encode a dim object to a signal with greater energy. 2) They can only disperse, and not re-concentrate, energy. Every point spread function (under the constraints of non-negativity and linear shift invariance) can only map objects to signals that are blurrier versions of themselves.
@@
To find the best achievable distribution, we first use Information-Driven Encoder Analysis Learning (IDEAL) (See Encoder design via information maximization with IDEAL) to learn the optimal encoder. We then generate random objects and encode them with this optimal encoder to sample from the distribution of encoded signals, add noise, and estimate the mutual information. We use the PixelCNN-based estimator [^2] We used PixelCNN for convenience, though estimators designed for 1D signals [uriaNeuralAutoregressiveDistribution2016, germainMADEMaskedAutoencoder2015a] might provide better accuracy., treating the signals as 2D images to match the estimator’s design.
@@
Next, we tested the effect of signal bandwidth on information capacity. Optical imaging systems are often characterized in terms of their space-bandwidth product [parkReviewBioopticalImaging2021], and the space-bandwidth product is often used synonymously with the word “information.” A more accurate term for the space-bandwidth product is “degrees of freedom,” since it quantifies the potential complexity of an electromagnetic field wave propagating through the system [franceschettiWaveTheoryInformation2017]. Information (in Shannon’s entropy/mutual information sense) also depends on the signal-to-noise ratio and the object-dependent ability of encoders to map to distinct signals. In our numerical simulation, the spatial extent of signals is fixed, so increasing the bandwidth of the signal increases the space-bandwidth product. We found that captured information increases linearly with the space-bandwidth product, with rate of increase depending on the object distribution (Fig. S7b, center).
@@
Finally, we examined the effect of sampling density on the amount of information that can be encoded. For a fixed bandwidth, the sampling density determines the number of pixels in the signal. The Nyquist sampling theorem [nyquistCertainTopicsTelegraph1928, shannonCommunicationPresenceNoise1949] states that a bandlimited signal can be perfectly reconstructed from its samples if the sampling density is at least twice the bandwidth. However, in the presence of noise, even when sampling at the Nyquist rate, there remains residual uncertainty about the signal, and additional samples can reduce this uncertainty. Experimentally, we found that, as expected, increasing the sampling density increases the amount of information, even beyond the Nyquist rate (Fig. S7b, right). However, the additional increases in information were progressively smaller. These results were consistent across both different object distributions and different average signal-to-noise ratios (for a delta function object distribution).
@@
$\delta(\mathbf{o}-\mathbf{o}_{1})$ is $1$ when the object is $\mathbf{o}_{1}$ and 0 otherwise, and similarly for $\delta(\mathbf{o}-\mathbf{o}_{2})$.
@@
Since the goal of the imaging system in this simple example is to classify the object as being a single point source or two point sources [harrisResolvingPowerDecision1964], the decoder is a binary classifier that takes in the noisy measurement and outputs a decision as to whether the object was a single point or two points.
@@
Our basic framework works for many imaging systems, including many optical systems whose purpose is to capture the object in as much detail as possible. Here we present an expanded model that accounts for specialized cases involving stochastic encoders and task-specific information, with detailed treatments provided in subsequent sections.
@@
The minimal model assumes deterministic encoders where the same object always produces the same noiseless image. In specialized cases, this assumption may not hold due to quantum effects in image formation or classical system variations like mechanical drift and illumination fluctuations. We provide a theoretical framework for handling such stochastic encoders in Section S2.7.
Additionally, the minimal model treats all object information as equally valuable, but many imaging applications focus on specific tasks that may require only a subset of the total object information. We develop extensions for task-specific information estimation in Section S2.8.
@@
As described in Mathematical framework, our probabilistic model represents noiseless images as a random 2D array $\mathbf{X}$ with distribution $p(\mathbf{x})$, where each element quantifies the energy at a single pixel. The corresponding noisy measurements form array $\mathbf{Y}$ with distribution $p(\mathbf{y})$. These distributions determine how much object-relevant information survives the measurement process.
@@
While imaging system performance typically varies across the field of view, we can quantify overall information content using a single value: the mutual information rate (information per pixel). This parallels how optical resolution is often characterized by a single number despite spatial variations [goodmanIntroductionFourierOptics2005].
@@
Mutual information estimation is a well-studied problem in many fields, including machine learning, neuroscience, and computational biology, and a number of different approaches have been proposed, some of which attempt to estimate mutual information directly [kraskovEstimatingMutualInformation2004a, darbellayEstimationInformationAdaptive1999, rossMutualInformationDiscrete2014], and others which try to infer its value by computing upper or lower bounds [mcallesterFormalLimitationsMeasurement2018, pooleVariationalBoundsMutual2019]. It is a challenging statistical problem to solve in general, and many estimators and bounds are known to suffer from high bias and/or variance, particularly on high-dimensional problems [mcallesterFormalLimitationsMeasurement2018, pooleVariationalBoundsMutual2019].
@@
Like mutual information estimation, there is a large body of literature on entropy estimation and many different approaches have been proposed [beirlantNonparametricEntropyEstimation1997], including those that form an estimate of the probability density function $\hat{p}(\mathbf{y})$ and plug it into the definition of entropy and those that estimate entropy based on the similarity of samples from $p(\mathbf{y})$ [kozachenkoSampleEstimateEntropy1987, lombardiNonparametricKnearestNeighbour2016, jiaoNearestNeighborInformation2018]. Both approaches, however, face difficulties in high dimensions.
@@
(A technical note: in optical systems, the outcomes of the random variables $\mathbf{X}$ and $\mathbf{Y}$ in our model are discrete, because they are 2D arrays of pixels, where each pixel takes on an intensity value that is an integer number of photons. However, for computational simplicity, we will approximate them as continuous random variables and use differential entropy instead of discrete entropy [coverElementsInformationTheory2005, mackayInformationTheoryInference2003]. These approximations break down at very low photon counts, less than $\sim 20$ photons, so in this paper, we use only data with photon counts greater than this. A possible direction of future work is to extend our approach to work with discrete random variables, which would allow it to be applied to data with lower photon counts.)
@@
$p(\mathbf{y}\mid\mathbf{x})$ embodies the various sources of noise in the detection process, including photon shot noise, detector read noise, etc. Here, we utilize established analytic models of detection noise in optical imaging [goodmanStatisticalOptics2015]. Empirical results suggest that the true noise in systems deviates from these models in low-light conditions [monakhovaDancingStarsVideo2022], but since our experiments are conducted in the high-light regime, we will assume that these models are accurate for the purposes of this paper. A possible direction for future work is to learn more accurate noise models from the data, as was done in [krullProbabilisticNoise2VoidUnsupervised2019].
@@
The entropy of a (scalar) Gaussian distribution $\mathcal{N}(\mu,\sigma)$ is [coverElementsInformationTheory2005]:
@@
Plugging this result into Equations 16 and 12 yields:
@@
Images with high photon counts are fundamentally limited by shot noise - randomness in photon arrival times due to the quantum nature of light. This shot noise follows a Poisson distribution with rate parameter equal to the expected number of photons at each pixel. When shot noise is the dominant source of noise, it can be accurately approximated by a Gaussian distribution with equal mean and variance [goodmanStatisticalOptics2015].
@@
We estimate this entropy by computing an upper bound, an approach that has proven more accurate than alternative bounds in high-dimensional settings [mcallesterFormalLimitationsMeasurement2018]. Our method fits a parametric model $p_{\mathbf{\theta}}(\mathbf{y})$ to the empirical distribution of noisy images using maximum likelihood estimation. The optimal parameters $\hat{\theta}_{\mathrm{MLE}}$ are found by minimizing the negative log likelihood of the observed data:
@@
This loss function, $\mathbb{E}\left[\log p_{\theta}(\mathbf{Y})\right]$, is also known as the cross-entropy between the model distribution $p_{\theta}(\mathbf{y})$ and the empirical distribution $p(\mathbf{y})$.
@@
Other models could be used within this framework, provided they allow direct evaluation of likelihood functions. Recent neural network architectures such as transformers [vaswaniAttentionAllYou2017], normalizing flows [kobyzevNormalizingFlowsIntroduction2021], and diffusion models [sohl-dicksteinDeepUnsupervisedLearning2015, songHowTrainYour2021] offer promising directions for future work.
While generative models are often evaluated by the visual quality of their samples, this metric does not necessarily correlate with cross-entropy performance [theisNoteEvaluationGenerative2016] - though some state-of-the-art models can achieve both [kingmaVariationalDiffusionModels]. We selected PixelCNN for its demonstrated effectiveness in minimizing test set cross-entropy [salimansPixelCNNImproving2017], combined with its relative simplicity and computational efficiency. However, the rapid advancement of generative models suggests that more efficient architectures providing tighter entropy bounds for a given computational efficiency are likely to emerge.
@@
To reduce model complexity and quantify performance over the full field of view with a single scalar quantity, we can make the simplifying assumption that images are stationary stochastic processes. This is the probabilistic analog of assuming constant optical resolution across the field of view [goodmanIntroductionFourierOptics2005].
@@
Stationary processes have another useful property: a constant entropy per pixel, called the entropy rate [coverElementsInformationTheory2005] (explained in Figure 8 of [pinkardVisualIntroductionInformation2022]). This rate can be defined in two equivalent ways:
@@
Our most flexible model is the PixelCNN [oordPixelRecurrentNeural2016, vandenoordConditionalImageGeneration2016, salimansPixelCNNImproving2017], which uses neural networks to capture complex pixel dependencies through an autoregressive approach.
@@
This factorization does not require specific assumptions about the joint distribution, so in theory it can be applied to any distribution. Creating the full model requires $D$ conditional distribution models, each taking between 1 and D1D-1 previous pixel values as input and outputting a 1D probability distribution for the next pixel. While this could be implemented with $D$ separate models, PixelCNN uses masked convolutions in a single neural network to model all conditionals simultaneously, dramatically reducing computational complexity.
Following previous work [oordPixelRecurrentNeural2016, vandenoordConditionalImageGeneration2016], our architecture uses a series of masked convolutions to maintain the autoregressive ordering, producing a conditional probability distribution at each pixel.
However, we make one crucial modification: instead of the standard softmax output layer, we use a mixture of Gaussians parameterized by the network (a mixture density network [bishopMixtureDensityNetworks1994]). This strategy is capable in theory of approximating any conditional probability distribution, in the same way that a neural network is capable of approximating any function [hornikMultilayerFeedforwardNetworks1989]. This modification is essential for estimating entropy in our framework, because it means the output distributions will be continuous probability densities, instead of discrete probability mass functions. Since our noise models $p(\mathbf{y}\mid\mathbf{x})$ are continuous (Gaussian), our entropy estimates for $H(\mathbf{Y})$ must also use continuous distributions to allow proper subtraction of the conditional entropy $H(\mathbf{Y}\mid\mathbf{X})$.
@@
While our models’ primary purpose is likelihood evaluation for entropy estimation (Section S2.4), generating samples provides valuable insight into model behavior. For simple models like the stationary Gaussian process, sample quality correlates well with likelihood accuracy, making visual inspection a useful diagnostic tool. However, this correlation breaks down for more complex models like PixelCNN [theisNoteEvaluationGenerative2016], though samples remain helpful for understanding model behavior.
@@
Computing the likelihood over an image larger than the length $D$ model the pixel was trained on can be accomplished by adding additional terms to this sum. For example, computing the log likelihood of an $D+1$ length image would require adding an additional term to the sum of the log likelihood of the the $D+1$ pixel conditioned on the previous D1D-1 pixels (the maximum extent of the model) that preceded it:
@@
Details of the likelihood computations for PixelCNNs are described in [oordPixelRecurrentNeural2016]. For stationary Gaussian processes, there is a closed form solution for finding the mean $\mu_{D}$ and variance $\sigma^{2}_{D}$ of a 1-dimensional Gaussian distribution conditioned on D1D-1 previous values. This involves decomposing the covariance matrix into a top left $(D-1)\times(D-1)$ block $\mathbf{\Sigma}_{1,1}$, a top right $(D-1)\times 1$ column vector $\mathbf{\Sigma}_{1,2}$, and a bottom left $1\times(D-1)$ row vector $\mathbf{\Sigma}_{2,1}$:
$\mathbf{\Sigma}=\begin{bmatrix}\mathbf{\Sigma}_{1,1}&\mathbf{\Sigma}_{1,2}\\ \mathbf{\Sigma}_{2,1}&\sigma^{2}_{M}\\ \end{bmatrix}$ (43)
Given D1D-1 previous values $\mathbf{y}_{D-1}$, the mean and variance of the $D$th value can be computed as:
@@
Similarly, sampling images larger than the patch size on which the models were trained can be accomplished by iteratively sampling each pixel conditioned on the previous D1D-1 pixels.
S2.7 Stochastic encoders
The deterministic encoder assumption underlying our framework breaks down in two scenarios:
First, in quantum imaging regimes, image formation becomes inherently random and cannot be predicted from Maxwell’s equations. These are beyond the scope of our present work.
Second, in systems with encoder uncertainty–deterministic but unpredictable variations like illumination fluctuations or mechanical drift that cause the same object to produce different noiseless images across measurements.